OpenStudy (anonymous):

What are the three means between 10 and 2,560? Select one: a. 40, 160, 640 b. -40, 160, 640 c. ±40, ±160, ±640 d. None of the above

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (gorv):

which mean you are talking about arithmatic geomatric or ......????

OpenStudy (gorv):

generally mean of a & b = (a+b)/2

OpenStudy (johnweldon1993):

Geometric means based on your last questions? If so, lets use the formula \[\large tn = t_1r^{n - 1}\] \(\large t_n\) would be the 2560 \(\large t_1\) is the 10 you have \(\large n\) would be 5 \[\large 2560 = 10(r^{5 - 1})\] So lets divide by 10 \[\large 256 = r^4\] \[\large r = \pm 4\] We can use that to solve for the other means... start with 6 and multiply it by 4..then continue to multiply the results by 4 until you get to 2560 we will also have 2 sets of means here (notice the \(\large \pm4\)) so we also need to multiply the 6 by -4...then do the same multiplying the results you get by -4 until you get to 2560

OpenStudy (anonymous):

Would it be c....

OpenStudy (johnweldon1993):

sorry meant starting with 10... and multipying by 4 and -4 lol:P

OpenStudy (johnweldon1993):

so lets see 10 times 4 = 40..times 4 = 160 times 4 = 640 and 10 times -4 = -40 times -4 = 160 times -4 = -640 so the list will either look like \[\large 40 , 160 , 640\] or \[\large -40, 160, -640\]

OpenStudy (anonymous):

So it must be A then right? 40,160,640..

OpenStudy (johnweldon1993):

Correct, sorry about that lol